# Stochastic Bregman Proximal Gradient Method Revisited: Kernel Conditioning and Painless Variance Reduction

### v3

@@ 1 Introduction
$x_{k+1}=\mathbf{T}_{\phi,h}^{\lambda}\big{(}x_{k},\nabla f(x_{k})\big{)}\qquad\mbox{with}\qquad\mathbf{T}_{\phi,h}^{\lambda}(x,v):=\mathop{\mathrm{argmin}}_{y\in\mathbb{R}^{d}}\,\,y^{\top}v+\phi(y)+\lambda^{-1}D_{h}(y,x),\vspace{-0.1cm}$ (3)
@@
{mdframed}
[backgroundcolor=gray!20,leftmargin=15pt, rightmargin=15pt] Q-1. Is the mismatch factor $C_{\mathcal{M}}$ uniformly bounded over $\mathcal{P}$ for the popular nonstandard measures in current literature? If not, how to interpret the existing complexity results? Can we find a new measure with constant mismatch and analyze complexity under it? This question is closely related to the instance-dependent nature of existing works. Take the squared primal gradient mapping $\mathcal{M}(\cdot)=\|\mathcal{G}_{\phi,h}^{\lambda}(\cdot)\|^{2}$ for example, existing results output a point $\bar{x}$ such that $\|\mathcal{G}_{\phi,h}^{\lambda}(\bar{x})\|^{2}\leq\epsilon$ with $O(\epsilon^{-2})$ samples [8, 9, 45]. If the mismatch $C_{\mathcal{M}}$ is uniformly bounded for the output over all instances in $\mathcal{P}$, then existing results immediately translates to the same $O(\epsilon^{-2})$ sample complexity for finding some $\bar{x}$ such that $\mathrm{dist}^{2}(0,\partial\Psi(\bar{x}))\leq\epsilon$, except for an additional constant factor (mismatch upper bound) hidden in big-O. Such a result will be independent of the instances in $\mathcal{P}$, and is thus an instance-free (worst-case) complexity. Unfortunately, this does not hold.
@@
How to provably accelerate stochastic BPG?    Current framework for analyzing stochastic BPG methods is mostly the classic vanilla SA approach based on the generalized descent lemma [3] and the bounded variance assumption, see e.g. [8, 9, 45, 47, 14]. However, more recent acceleration techniques for stochastic first-order algorithms, like momentum [28], shuffling [31, 32], and variance reduction [38, 39, 44, 7], all share a same analysis workflow that bounds the stochastic errors by gradient differences, and then controls gradient differences by iterative descents through gradient Lipschitz property. Unfortunately, the smooth-adaptability condition alone is not able to ensure an appropriate Lipschitz-like bound for gradient differences. Due to this issue, even under the nonstandard (instance-dependent) stationarity measures, only the typical $O(\epsilon^{-2})$ complexity for vanilla SA approach was obtained. Therefore, the second question in this paper arises: {mdframed}[backgroundcolor=gray!20,leftmargin=15pt, rightmargin=15pt] Q-2. Is there a natural kernel regularity condition that is able to ensure an appropriate Lipschitz-like bound for gradient differences under smooth-adaptability? If such a Lipschitz-like bound can be established, then it may function as a general tool for analyzing momentum, shuffling or variance reduced stochastic BPG methods under smooth-adaptability.
@@
Notations. For $\forall x\in\mathbb{R}^{d}$, We default $\|x\|:=\sqrt{x^{\top}x}$ the $\ell_{2}$-norm, and we denote $\|x\|_{1}:=\sum_{i}|x_{i}|$ as the $\ell_{1}$-norm. For a matrix $X\in\mathbb{R}^{d\times d}$, we use $\lambda_{\max}(X)$ and $\lambda_{\min}(X)$ to denote the maximal and minimal eigenvalues of $X$, respectively. And we use $\|X\|$ to denote the $\ell_{2}$-operator norm of $X$. When $X$ is positive semidefinite, we write $X\succeq 0$, and we have $\|X\|=\lambda_{\max}(X)$ in this case. For any set $\mathcal{X}$, we denote $\mathrm{id}_{\mathcal{X}}(\cdot)$ as the indicator function of the set. Namely, $\mathrm{id}_{\mathcal{X}}(x)=0$ if $x\in\mathcal{X}$ and $\mathrm{id}_{\mathcal{X}}(x)=+\infty$ if $x\notin\mathcal{X}$. We denote the interior of $\mathcal{X}$ as $\mathrm{int}(\mathcal{X})$ and we denote the boundary of $\mathcal{X}$ as $\partial\mathcal{X}$. We denote $[n]:=\{1,2,\cdots,n\}$. For any vector $x\in\mathbb{R}^{d}$, the function $\mathrm{sign}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ returns a sign vector of $x$. That is, for any $s=\mathrm{sign}(x)$, its $i$-th element satisfies $s_{i}=1$ if $x_{i}\geq 0$ and $s_{i}=-1$ if $x_{i}<0$, for any $1\leq i\leq d$. Because many literature use the terminology $L$-smooth to denote $L$-Lipschitz continuity of the gradient, we will use “continuously differentiable” instead of “smooth” to avoid confusion.
@@ 2.1 Preliminary results
Let $f$ and $h$ be twice continuously differentiable in $\mathbb{R}^{d}$, and let $h$ be strictly convex. Then we assume $f$ is $L$-smooth adaptable to $h$ for some positive constant $L>0$. In other words, both $Lh+f$ and $Lh-f$ are convex functions.
@@
$-L\nabla^{2}h(x)\preceq\nabla^{2}f(x)\preceq L\nabla^{2}h(x),\,\,\forall x\in\mathbb{R}^{d}.$ (6)
[…]
@@ 2.2 Instance-dependent nature of existing complexity results
$\|\mathcal{G}_{\phi,h}^{\lambda}(x)\|^{2}\leq{\color[rgb]{0,0,0}\frac{2}{\mu}B_{\lambda}(x)}\qquad\mbox{and}\qquad\|\mathcal{G}_{\phi,h}^{\lambda}(x)\|^{2}\leq\frac{4}{\mu}M_{\lambda}(x).$
@@
The set of Fréchet subgradient of $\Psi$ at $x$ is called the Fréchet subdifferential and is denoted as ^$\hat{\partial}\Psi(x)$. Then the limiting Fréchet subdifferential denoted by $\partial\Psi(x)$ is defined as
$\partial\Psi(x)=\{v:\mathrm{there\leavevmode\nobreak\ }\exists\leavevmode\nobreak\ x_{k}\rightarrow x\leavevmode\nobreak\ \mathrm{and}\leavevmode\nobreak\ v_{k}\in\hat{\partial}\Psi(x_{k})\leavevmode\nobreak\ \mathrm{s.t.}\leavevmode\nobreak\ v_{k}\rightarrow v\}.$
@@
For any $x,v\in\mathbb{R}^{d}$ and any strictly convex kernel $h$, denote $x_{\lambda}^{+}(v):=\mathbf{T}_{\phi,h}^{\lambda}(x,v)$. Let $[x,x_{\lambda}^{+}(v)]$ be the line segment between $x_{\lambda}^{+}(v)$ and $x$, then there exists $u\in\partial\phi(x_{\lambda}^{+}(v))$ such that
@@
${\color[rgb]{0,0,0}\frac{\mathrm{dist}^{2}(0,\partial\Psi(x_{\lambda}^{+}))}{B_{\lambda}(x)}\leq 8\kappa_{h}(\mathcal{X})L_{h}(\mathcal{X})},\qquad\quad\frac{\mathrm{dist}^{2}(0,\partial\Psi(x_{\lambda}^{+}))}{M_{\lambda}(x)}\leq 4\kappa_{h}(\mathcal{X})L_{h}(\mathcal{X}),$
@@
${\color[rgb]{0,0,0}\frac{\|\nabla\Psi(x)\|^{2}}{B_{\lambda}(x)}\leq 2\kappa_{h}(\mathcal{X})L_{h}(\mathcal{X})},\qquad\frac{\|\nabla\Psi(x)\|^{2}}{M_{\lambda}(x)}\leq\kappa_{h}(\mathcal{X})L_{h}(\mathcal{X}),\quad\mbox{and}\quad\,\,\frac{\|\nabla\Psi(x)\|^{2}}{\|\mathcal{G}_{\phi,h}^{\lambda}(x)\|^{2}}\leq L_{h}^{2}(\mathcal{X}).$
As discussed in the introduction, the mismatch factors in Corollary 2.6 stand for the ability for a stationarity measure to upper bound the the Fréchet measure. Take the case $\phi=0$ for example, with $\mathcal{X}=[x,x_{\lambda}^{+}]$, Corollary 2.6 indicates that having $\|\mathcal{G}_{\phi,h}^{\lambda}(x)\|^{2}\leq\epsilon$ only implies $\|\nabla\Psi(x)\|^{2}\leq L_{h}^{2}(\mathcal{X})\epsilon$. Therefore, smaller and uniformly bounded mismatch factors are always more desirable, at least for the output. However, for general smooth-adaptable setting, the $L_{h}(\mathcal{X})$ factor in Corollary 2.6 can be unbounded over $\mathbb{R}^{d}$ for many popular non-Lipschitz-smooth kernels. Depending on the kernel Hessian around the output, if the instance is simple and the $L_{h}(\mathcal{X})$ factor is mild, then the existing $O(\epsilon^{-1})$ complexity in terms of $B_{\lambda}(\cdot)$, $M_{\lambda}(\cdot)$, or $\|\mathcal{G}_{\phi,h}^{\lambda}(\cdot)\|^{2}$ can be directly translated to that of the standard Fréchet measure. On the other hand, for hard instances where $L_{h}(\mathcal{X})$ is huge or even unbounded, these results will fail to provide meaningful complexity for Fréchet measure or finite-step predictions for solution quality. The same instance-dependence happens for the stochastic setting, where the only difference is that for stochastic algorithms whose theoretical complexity is often established for a randomly selected output from for all iterations, the mismatch should also consider all iterations.
@@
It can be observed in the first figure that even when the Bregman residuals and primal gradient mapping are small, the Fréchet measure, which reduces to gradient size when $\phi=0$, still remains large. For the mismatch factor discussed in Corollary 2.6, take the squared primal gradient mapping for example, it grows to around $10^{8}$ while still not reaching the limit, which illustrates the mismatch issue for hard instances in our discussion. For the second and third figures, we illustrate an interesting “rate inversion” phenomenon, where we apply polynomial kernels of different degree $r$ to the same problem and report the convergence rate w.r.t. Fréchet measure and primal gradient mapping, respectively. From the second figure, it is suggested that higher degree kernels converges “faster” in terms of primal gradient mapping. However, if we look at Fréchet measure in the third figure, the quantity that we truly care about, kernels with lower degrees are more appropriate. Though we only plot primal gradient mapping, the same phenomenon also happens for the two Bregman residual measures. For the last figure, we plot $\log_{2}\|\nabla\Psi(x_{k})\|^{2}$ versus $\log_{2}k$, whose slope represents the exponent of an $O(T^{-\gamma})$ sublinear rate. From the experiment, it is clear that the polynomial kernels of different degree $r$ result in different exponent $\gamma$ in the sublinear convergence rate, while none of them exhibits a $-1$ slope that corresponds to an $O(1/T)$ rate in the existing BPG literature. This illustrates the inability of instance-dependent bounds to characterize hard instances, while the actual worst-case complexity or convergence rate still remains unknown for BPG methods.
@@ 2.4 The kernel conditioning regularity
Basically, Assumption 2.9 states that when the compact set $\mathcal{X}$ is not very large, the localized condition number $\kappa_{h}(\mathcal{X})$ over $\mathcal{X}$ will be uniformly bounded, even when both $\lambda_{\min}(\nabla^{2}h(x))$ and $\lambda_{\max}(\nabla^{2}h(x))$ go to +$\infty$+ ∞. To the best of our knowledge, such a kernel regularity condition has not been considered in the existing works. In later discussion, if a kernel $h$ satisfies Assumption 2.9, then we will say $h$ is KC-regular. And we will abbreviate kernel conditioning regularity as KC-regularity. As a result, Lemma 2.8 immediately indicates that when $\|x-x_{\lambda}^{+}\|\leq\delta$ for some predetermined $\delta>0$, the mismatch ${\mathrm{dist}^{2}(0,\partial\Psi(x_{\lambda}^{+}))}/{\|\mathcal{D}_{\phi,h}^{\lambda}(x)\|^{2}}\leq(1+L\lambda\kappa_{h}^{\delta})^{2}$ is at most a constant. In addition to KC-regularity, we also inherits the commonly adopted global $\mu$-strong convexity regularity assumption in this paper, from the existing works [5, 8, 9, 15, 16, 19, 26, 27, 33, 40, 47, etc.].
@@
$\kappa_{h+g}(\mathcal{X})$ ::=:= $\frac{\sup\{\lambda_{\max}(\nabla^{2}h(x)+\nabla^{2}g(x)):x\in\mathcal{X}\}}{\inf\{\lambda_{\min}(\nabla^{2}h(x)+\nabla^{2}g(x)):x^{\prime}\in\mathcal{X}\}}$
@@
Hence we complete the proof of Theorem 2.11.
@@
Let $h(x)\!=\!\frac{{\color[rgb]{0,0,0}\alpha}}{2}\|x\|^{2}+\frac{1}{r+2}\|x\|^{r+2}$ be a power kernel for some real number $r\geq 0$ and $\alpha>0$. Then this kernel satisfies the following properties:
@@
$\mathrm{diam}(\mathcal{X})\leq\frac{1}{r}\cdot\max\Big{\{}{\color[rgb]{0,0,0}\alpha^{1/r}}\,,\,\min_{u\in\mathcal{X}}\|u\|\Big{\}}.$

### v4

@@ 1 Introduction
$x_{k+1}=\mathbf{T}_{\phi,h}^{\lambda}\big(x_{k},\nabla f(x_{k})\big)\qquad\mbox{with}\qquad\mathbf{T}_{\phi,h}^{\lambda}(x,v):=\mathop{\mathrm{argmin}}_{y\in\mathbb{R}^{d}}\,\,y^{\top}v+\phi(y)+\lambda^{-1}D_{h}(y,x),\vskip-2.84544pt$ (3)
@@
Q-1. Is the mismatch factor $C_{\mathcal{M}}$ uniformly bounded over $\mathcal{P}$ for the popular nonstandard measures in current literature? If not, how to interpret the existing complexity results? Can we find a new measure with constant mismatch and analyze complexity under it?
This question is closely related to the instance-dependent nature of existing works. Take the squared primal gradient mapping $\mathcal{M}(\cdot)=\|\mathcal{G}_{\phi,h}^{\lambda}(\cdot)\|^{2}$ for example, existing results output a point $\bar{x}$ such that $\|\mathcal{G}_{\phi,h}^{\lambda}(\bar{x})\|^{2}\leq\epsilon$ with $O(\epsilon^{-2})$ samples [8, 9, 45]. If the mismatch $C_{\mathcal{M}}$ is uniformly bounded for the output over all instances in $\mathcal{P}$, then existing results immediately translates to the same $O(\epsilon^{-2})$ sample complexity for finding some $\bar{x}$ such that $\mathrm{dist}^{2}(0,\partial\Psi(\bar{x}))\leq\epsilon$, except for an additional constant factor (mismatch upper bound) hidden in big-O. Such a result will be independent of the instances in $\mathcal{P}$, and is thus an instance-free (worst-case) complexity. Unfortunately, this does not hold.
@@
How to provably accelerate stochastic BPG?    Current framework for analyzing stochastic BPG methods is mostly the classic vanilla SA approach based on the generalized descent lemma [3] and the bounded variance assumption, see e.g. [8, 9, 45, 47, 14]. However, more recent acceleration techniques for stochastic first-order algorithms, like momentum [28], shuffling [31, 32], and variance reduction [38, 39, 44, 7], all share a same analysis workflow that bounds the stochastic errors by gradient differences, and then controls gradient differences by iterative descents through gradient Lipschitz property. Unfortunately, the smooth-adaptability condition alone is not able to ensure an appropriate Lipschitz-like bound for gradient differences. Due to this issue, even under the nonstandard (instance-dependent) stationarity measures, only the typical $O(\epsilon^{-2})$ complexity for vanilla SA approach was obtained. Therefore, the second question in this paper arises:
Q-2. Is there a natural kernel regularity condition that is able to ensure an appropriate Lipschitz-like bound for gradient differences under smooth-adaptability?
If such a Lipschitz-like bound can be established, then it may function as a general tool for analyzing momentum, shuffling or variance reduced stochastic BPG methods under smooth-adaptability.
@@
Notations. For $\forall x\in\mathbb{R}^{d}$, We default $\|x\|:=\sqrt{x^{\top}x}$ the $\ell_{2}$-norm, and we denote $\|x\|_{1}:=\sum_{i}|x_{i}|$ as the $\ell_{1}$-norm. For a matrix $X\in\mathbb{R}^{d\times d}$, we use $\lambda_{\max}(X)$ and $\lambda_{\min}(X)$ to denote the maximal and minimal eigenvalues of $X$, respectively. And we use $\|X\|$ to denote the $\ell_{2}$-operator norm of $X$. When $X$ is positive semidefinite, we write $X\succeq 0$, and we have $\|X\|=\lambda_{\max}(X)$ in this case. For any set $\mathcal{X}$, we denote $\mathrm{id}_{\mathcal{X}}(\cdot)$ as the indicator function of the set. Namely, $\mathrm{id}_{\mathcal{X}}(x)=0$ if $x\in\mathcal{X}$ and $\mathrm{id}_{\mathcal{X}}(x)=+\infty$ if $x\notin\mathcal{X}$. We denote the interior of $\mathcal{X}$ as $\mathrm{int}(\mathcal{X})$ and we denote the boundary of $\mathcal{X}$ as $\partial\mathcal{X}$. We denote $n]:=\{1,2,\cdots,n\}$. For any vector $x\in\mathbb{R}^{d}$, the function $\mathrm{sign}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ returns a sign vector of $x$. That is, for any $s=\mathrm{sign}(x)$, its $i$-th element satisfies $s_{i}=1$ if $x_{i}\geq 0$ and $s_{i}=-1$ if $x_{i}<0$, for any $1\leq i\leq d$. Because many literature use the terminology $L$-smooth to denote $L$-Lipschitz continuity of the gradient, we will use “continuously differentiable” instead of “smooth” to avoid confusion.
@@ 2.1 Preliminary results
Let $f$ and $h$ be twice continuously differentiable in $\mathbb{R}^{d}$, and let $h$ be strictly convex. Then we assume $f$ is $L$-smooth adaptable to $h$ for some positive constant $L>0$. In other words, both $Lh+f$ and LhfLh-f are convex functions.
@@
$L\nabla^{2}h(x)\preceq\nabla^{2}f(x)\preceq L\nabla^{2}h(x),\,\,\forall x\in\mathbb{R}^{d}.$ (6)
As we consider the problem class where $\nabla f$ is not globally Lipschitz continuous, then naturally, one would expect $\|\nabla^{2}f(x)\|$ and $\|\nabla^{2}h(x)\|$ to grow unbounded in a $\limsup$ sense as $\|x\|\to+\infty$. A particularly interesting example that satisfies the smooth adaptability assumption is the function class with polynomially growing Hessian, as described below.
@@ 2.2 Instance-dependent nature of existing complexity results
$\|\mathcal{G}_{\phi,h}^{\lambda}(x)\|^{2}\leq\frac{2}{\mu}B_{\lambda}(x)\qquad\mbox{and}\qquad\|\mathcal{G}_{\phi,h}^{\lambda}(x)\|^{2}\leq\frac{4}{\mu}M_{\lambda}(x).$
@@
The set of Fréchet subgradient of $\Psi$ at $x$ is called the Fréchet subdifferential and is denoted as $\hat{\partial}\Psi(x)$. Then the limiting Fréchet subdifferential denoted by $\partial\Psi(x)$ is defined as
$\partial\Psi(x)=\{v:\mathrm{there\penalty 10000\ }\exists\penalty 10000\ x_{k}\rightarrow x\penalty 10000\ \mathrm{and}\penalty 10000\ v_{k}\in\hat{\partial}\Psi(x_{k})\penalty 10000\ \mathrm{s.t.}\penalty 10000\ v_{k}\rightarrow v\}.$
@@
For any $x,v\in\mathbb{R}^{d}$ and any strictly convex kernel $h$, denote $x_{\lambda}^{+}(v):=\mathbf{T}_{\phi,h}^{\lambda}(x,v)$. Let $x,x_{\lambda}^{+}(v)]$ be the line segment between $x_{\lambda}^{+}(v)$ and $x$, then there exists $u\in\partial\phi(x_{\lambda}^{+}(v))$ such that
@@
$\frac{\mathrm{dist}^{2}(0,\partial\Psi(x_{\lambda}^{+}))}{B_{\lambda}(x)}\leq 8\kappa_{h}(\mathcal{X})L_{h}(\mathcal{X}),\qquad\quad\frac{\mathrm{dist}^{2}(0,\partial\Psi(x_{\lambda}^{+}))}{M_{\lambda}(x)}\leq 4\kappa_{h}(\mathcal{X})L_{h}(\mathcal{X}),$
@@
$\frac{\|\nabla\Psi(x)\|^{2}}{B_{\lambda}(x)}\leq 2\kappa_{h}(\mathcal{X})L_{h}(\mathcal{X}),\qquad\frac{\|\nabla\Psi(x)\|^{2}}{M_{\lambda}(x)}\leq\kappa_{h}(\mathcal{X})L_{h}(\mathcal{X}),\quad\mbox{and}\quad\,\,\frac{\|\nabla\Psi(x)\|^{2}}{\|\mathcal{G}_{\phi,h}^{\lambda}(x)\|^{2}}\leq L_{h}^{2}(\mathcal{X}).$
As discussed in the introduction, the mismatch factors in Corollary 2.6 stand for the ability for a stationarity measure to upper bound the Fréchet measure. Take the case $\phi=0$ for example, with $\mathcal{X}=[x,x_{\lambda}^{+}]$, Corollary 2.6 indicates that having $\|\mathcal{G}_{\phi,h}^{\lambda}(x)\|^{2}\leq\epsilon$ only implies $\|\nabla\Psi(x)\|^{2}\leq L_{h}^{2}(\mathcal{X})\epsilon$. Therefore, smaller and uniformly bounded mismatch factors are always more desirable, at least for the output. However, for general smooth-adaptable setting, the $L_{h}(\mathcal{X})$ factor in Corollary 2.6 can be unbounded over $\mathbb{R}^{d}$ for many popular non-Lipschitz-smooth kernels. Depending on the kernel Hessian around the output, if the instance is simple and the $L_{h}(\mathcal{X})$ factor is mild, then the existing $O(\epsilon^{-1})$ complexity in terms of $B_{\lambda}(\cdot)$, $M_{\lambda}(\cdot)$, or $\|\mathcal{G}_{\phi,h}^{\lambda}(\cdot)\|^{2}$ can be directly translated to that of the standard Fréchet measure. On the other hand, for hard instances where $L_{h}(\mathcal{X})$ is huge or even unbounded, these results will fail to provide meaningful complexity for Fréchet measure or finite-step predictions for solution quality. The same instance-dependence happens for the stochastic setting, where the only difference is that for stochastic algorithms whose theoretical complexity is often established for a randomly selected output from for all iterations, the mismatch should also consider all iterations.
@@
It can be observed in the first figure that even when the Bregman residuals and primal gradient mapping are small, the Fréchet measure, which reduces to gradient size when $\phi=0$, still remains large. For the mismatch factor discussed in Corollary 2.6, take the squared primal gradient mapping for example, it grows to around $10^{8}$ while still not reaching the limit, which illustrates the mismatch issue for hard instances in our discussion. For the second and third figures, we illustrate an interesting “rate inversion” phenomenon, where we apply polynomial kernels of different degree $r$ to the same problem and report the convergence rate w.r.t. Fréchet measure and primal gradient mapping, respectively. From the second figure, it is suggested that higher degree kernels converges “faster” in terms of primal gradient mapping. However, if we look at Fréchet measure in the third figure, the quantity that we truly care about, kernels with lower degrees are more appropriate. Though we only plot primal gradient mapping, the same phenomenon also happens for the two Bregman residual measures. For the last figure, we plot $\log_{2}\|\nabla\Psi(x_{k})\|^{2}$ versus $\log_{2}k$, whose slope represents the exponent of an $O(T^{-\gamma})$ sublinear rate. From the experiment, it is clear that the polynomial kernels of different degree $r$ result in different exponent $\gamma$ in the sublinear convergence rate, while none of them exhibits a 1-1 slope that corresponds to an $O(1/T)$ rate in the existing BPG literature. This illustrates the inability of instance-dependent bounds to characterize hard instances, while the actual worst-case complexity or convergence rate still remains unknown for BPG methods.
@@ 2.4 The kernel conditioning regularity
Basically, Assumption 2.9 states that when the compact set $\mathcal{X}$ is not very large, the localized condition number $\kappa_{h}(\mathcal{X})$ over $\mathcal{X}$ will be uniformly bounded, even when both $\lambda_{\min}(\nabla^{2}h(x))$ and $\lambda_{\max}(\nabla^{2}h(x))$ go to $\infty$. To the best of our knowledge, such a kernel regularity condition has not been considered in the existing works. In later discussion, if a kernel $h$ satisfies Assumption 2.9, then we will say $h$ is KC-regular. And we will abbreviate kernel conditioning regularity as KC-regularity. As a result, Lemma 2.8 immediately indicates that when $\|x-x_{\lambda}^{+}\|\leq\delta$ for some predetermined $\delta>0$, the mismatch ${\mathrm{dist}^{2}(0,\partial\Psi(x_{\lambda}^{+}))}/{\|\mathcal{D}_{\phi,h}^{\lambda}(x)\|^{2}}\leq(1+L\lambda\kappa_{h}^{\delta})^{2}$ is at most a constant. In addition to KC-regularity, we also inherits the commonly adopted global $\mu$-strong convexity regularity assumption in this paper, from the existing works [5, 8, 9, 15, 16, 19, 26, 27, 33, 40, 47, etc.].
@@
$\kappa_{h+g}(\mathcal{X})$ $:=$ $\frac{\sup\{\lambda_{\max}(\nabla^{2}h(x)+\nabla^{2}g(x)):x\in\mathcal{X}\}}{\inf\{\lambda_{\min}(\nabla^{2}h(x)+\nabla^{2}g(x)):x^{\prime}\in\mathcal{X}\}}$
@@
Hence we complete the proof of Theorem 2.11. ∎
@@
Let $h(x)=\frac{\alpha}{2}\|x\|^{2}+\frac{1}{r+2}\|x\|^{r+2}$ be a power kernel for some real number $r\geq 0$ and $\alpha>0$. Then this kernel satisfies the following properties:
@@
$\mathrm{diam}(\mathcal{X})\leq\frac{1}{r}\cdot\max\Big\{\alpha^{1/r}\,,\,\min_{u\in\mathcal{X}}\|u\|\Big\}.$